The ABC fund, a close end fund, consist of three securities- 1000 shares of Security A, which is currently trading at $35; 2000 shares of Security B, which is trading at $45; and 3000 shares of Security C, which is trading at $55;  

What is ABC funds' NAV if has $87 000 in net liabilities and 10 000 shares outstanding?
Answer #1
N P V=N*P
Number of Shares Current Market Price Value
SECURITY A 1000 $35 $35,000
SECURITY B 2000 $45 $90,000
SECURITY C 3000 $55 $165,000
Total $290,000
TV Total Value of securities $290,000
L Net Liability $87,000
AV=TV-L Asset Value $203,000
n Number of shares outstanding 10000
NAV=AV/n Net Asset Value (NAV) $20.30
NAV=$20.30